• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Chao

    Log entry time 16:06:59 on March10,2012

    Entry number 365421

    keyword=Day Shift Summary

    Shift Workers:
    C. Gu (SL), Y. Zhao (TO), Y. Han (3rd), with optics, beamline and target experts

    08:00 Shift starts, Jixie is doing beam position scan for optics

    08:03 move beam to (0,4)

    09:00 move beam to (0,-4) but it never settled

    09:05 Beam is down, asked for a controlled access to let Jack get in to fix the right dipole

    11:30 After discussed with Alexandre, we called the target group and they start to prepare for doing TE

    12:50 Jack ramped the right dipole down and reset it. The magnet seems to behave normally, start to ramp it up for Optics

    Meanwhile Alexandre is doing BCM Calibration

    15:00 it is decided we will do GEP production first, so we start to set the HRS momentum to 2.228GeV and septa current to 718.1A

    16:00 Left arm is locked and right dipole is still moving